# WildScore: Benchmarking MLLMs in-the-Wild Symbolic Music Reasoning

### v1

@@ Abstract
Recent advances in Multimodal Large Language Models (MLLMs) have demonstrated impressive capabilities across various vision-language tasks. However, their reasoning abilities in the multimodal symbolic music domain remain largely unexplored. We introduce WildScore, the first in-the-wild multimodal symbolic music reasoning and analysis benchmark, designed to evaluate MLLMs’ capacity to interpret real-world music scores and answer complex musicological queries. Each instance in WildScore is sourced from genuine musical compositions and accompanied by authentic user-generated questions and discussions, capturing the intricacies of practical music analysis. To facilitate systematic evaluation, we propose a systematic taxonomy, comprising both high-level and fine-grained musicological ontologies. Furthermore, we frame complex music reasoning as multiple-choice question answering, enabling controlled and scalable assessment of MLLMs’ symbolic music understanding. Empirical benchmarking of state-of-the-art MLLMs on WildScore reveals intriguing patterns in their visual-symbolic reasoning, uncovering both promising directions and persistent challenges for MLLMs in symbolic music reasoning and analysis. We release the dataset111 https://huggingface.co/datasets/GM77/WildScore and code [^2] https://github.com/GaganVM/WildScore .
@@ 1 Introduction
In this work, we present WildScore, the first multimodal symbolic music reasoning benchmark constructed from in-the-wild data. WildScore comprises real music scores by actual composers, paired with user-generated questions and discussions sourced from public forums. Many real-world queries require integrating several musical reasoning steps, including identifying notational symbols, interpreting harmonic progressions, and contextualizing expressive markings, which demand the need for MLLMs that can perform compositional and context-aware multimodal reasoning. This collection reflects the authentic diversity and complexity of symbolic music interpretation as it occurs in real-world discourse, and demands nuanced reasoning about notation, structure, and musical intent Xu et al. (2024) .
@@
MusicTheory-
Bench
Czajka et al. (2024b)
Expert-curated Theory MCQ 2 Manual Acc.
MAESTRO
Hawthorne et al. (2019)
Curated
(Competition recs.)
Audio 2 Automatic F1-score
MusicNet
Thickstun et al. (2017b)
Curated
(Classical recs.)
Audio N/A Manual Recall
NES-MDB
Donahue et al. (2018)
Curated
(Game Audio)
Audio 11 Automatic N/A
MusicScore
Chou et al. (2023)
Curated
(Public Scores)
Score image corpus 8 Automatic FID
Lakh MIDI
Raffel (2016)
In-the-wild
(Web-MIDI)
MIDI N/A Automatic N/A
@@ 2.1 Symbolic Music Understanding and Benchmarks
Symbolic music understanding has traditionally been evaluated using clean, structured datasets such as MusicNet (Thickstun et al., 2017a), NES-MDB (Donahue et al., 2018), and MAESTRO (Hawthorne et al., 2019). These datasets align audio with symbolic formats to facilitate tasks like transcription and generation. However, they reflect highly curated environments, lacking the variability, ambiguity, and informal nature of user-generated content. Other symbolic corpora like MusicScore Chou et al. (2023) or Lakh MIDI Raffel (2016) further extend coverage but remain either score-centric or MIDI-based without real-world contextual grounding. Recent efforts like MusicTheoryBench Czajka et al. (2024b) introduce theory-centric evaluations, but they rely on expert-curated questions in controlled settings. WildScore differs by grounding symbolic music analysis in online discourse, incorporating informal reasoning and context-dependent ambiguity from platforms like Reddit Reddit (2024).
@@ 3 WildScore
The aim of this study is to evaluate the visual context understanding of Multimoda Large Language Models (MLLMs) for symbolic musical score as shown in Figure 2. To this end, we introduce WildScore. In this section, we describe the dataset details; Section 4 presents the evaluation of vision–language reasoning over symbolic musical scores.
Our dataset creation process involves two distinct phases: (1) data collection (3.1), (2) multimodal filtering (3.2). Together, these phases enforce sample relevance, symbol–image grounding, and rigorous quality control, yielding a benchmark that robustly evaluates MLLMs’ visual context understanding of symbolic music.
@@ 3.3 Dataset Categorization
To support structured analysis and evaluation, we categorized our dataset into five categories as shown in Figure  3 to represent core aspects of music theory. These categories are further divided into twelve detailed subcategories as shown in Figure  4:
@@ 3.4 Dataset Overview & Statistics
We assign each question to Easy, Medium, or Hard using an LLM-based rubric. Specifically, we prompt GPT-4.1 with a few-shot template designed from examples curated by a Level 5 expert (criteria in Table 6) to rate the expected difficulty from the prompt. The resulting distribution is as shown in Table 7. Additional construction details, annotator instructions, and prompt templates are provided in the appendix 5.
@@ 4.2 Quantitative Results
GPT-4.1-mini OpenAI (2023)
Params: undisclosed
w/ Image 72.12 69.57 70.14 63.20 64.15 68.31
@@
Qwen-2.5-VL Bai et al. (2023)
Params: 8.29B
w/ Image 52.88 52.17 47.06 47.20 58.49 49.73
@@
Phi-3-Vision Abdin et al. (2024)
Params: 4.15B
w/ Image 45.19 52.17 48.42 48.00 56.60 48.82
@@
Gemma-3 Team et al. (2025)
Params: 4.3B
w/ Image 40.27 52.24 47.89 43.55 53.42 46.34
@@
MiniCPM Hu et al. (2024)
Params: 3.43B
w/ Image 50.00 45.65 44.34 44.80 47.17 45.90
@@
InternVL Chen et al. (2023)
Params: 9.14B
w/ Image 46.15 36.96 36.65 37.60 43.40 39.34
@@
LLaVA Liu et al. (2023b)
Params: 7.06B
w/ Image 37.50 41.30 28.96 32.00 35.85 32.97
@@
Table 4: Per-subcategory accuracy (%) by model and input modality, with subcategories grouped by category.
A subcategory analysis (Table 4) reveals heterogeneous image contributions across models. For GPT-4.1-mini, accuracies peak on Dynamics and Articulation (87.18%) and Modal Mixture (79.25%) and drop on Orchestral Texture (33.33%) and Contrapuntal Forms (40.00%). Other systems show attenuated or even negative image gains in multiple subcategories. We hypothesize that this heterogeneity reflects differences in multimodal pretraining and alignment in models with stronger vision–language objectives and instruction tuning appear better grounded in symbolic notation, whereas those trained primarily on natural-image corpora or with weaker visual adapters show limited benefit from images. A Level-5 human expert was evaluated on 100 proportionally sampled questions spanning easy, medium, and hard categories, achieving an overall average accuracy of 72%. Additionally, besides Contrapuntal Forms most model performance without the image is better than random guessing, highlighting that the naturalized data used to create WildScore may not fully require perception of the scores. This makes an interesting contrast to recent synthetically difficult benchmarks that force multi-modal perception to succeed Zang et al. (2025), as such difficult benchmarks may not reflect the real distribution of questions, where perception may not always be necessary.
@@ 6 Conclusion
Empirical results indicate that while current state-of-the-art MLLMs exhibit substantial promise, particularly in tasks involving surface-level recognition and straightforward analysis, they continue to struggle significantly with deep symbolic abstraction, rhythmic complexity, and orchestration intricacies especially when presented as an image. Significant differences in performance demonstrated by popular mulimodal large language models between text-only and visual inputs highlight how important visual context is for precise musicological interpretation.

### v2

@@ Abstract
Recent advances in Multimodal Large Language Models (MLLMs) have demonstrated impressive capabilities across various vision-language tasks. However, their reasoning abilities in the multimodal symbolic music domain remain largely unexplored. We introduce WildScore, the first in-the-wild multimodal symbolic music reasoning and analysis benchmark, designed to evaluate MLLMs’ capacity to interpret real-world music scores and answer complex musicological queries. Each instance in WildScore is sourced from genuine musical compositions and accompanied by authentic user-generated questions and discussions, capturing the intricacies of practical music analysis. To facilitate a comprehensive evaluation, we propose a systematic taxonomy, comprising both high-level and fine-grained musicological ontologies. Furthermore, we frame complex music reasoning as multiple-choice question answering, enabling controlled and scalable assessment of MLLMs’ symbolic music understanding. Empirical benchmarking of state-of-the-art MLLMs on WildScore reveals intriguing patterns in their visual-symbolic reasoning, uncovering both promising directions and persistent challenges for MLLMs in symbolic music reasoning and analysis. We release the dataset [^1] https://huggingface.co/datasets/GM77/WildScore and code [^2] https://github.com/GaganVM/WildScore .
@@ 1 Introduction
In this work, we present WildScore, the first multimodal symbolic music reasoning benchmark constructed from in-the-wild data. WildScore comprises real music scores by actual composers, paired with user-generated questions and discussions sourced from public forums. Many real-world queries require integrating several musical reasoning steps, including identifying notational symbols, interpreting harmonic progressions, and contextualizing expressive markings, which demand the need for MLLMs that can perform compositional and context-aware multimodal reasoning. This collection reflects the authentic diversity and complexity of symbolic music interpretation as it occurs in real-world discourse, and demands nuanced reasoning about notation, structure, and musical intent Xu et al. (2024); Surana et al. (2022).
@@
MusicTheory- Bench Czajka et al. (2024b) Expert-curated Theory MCQ 2 Manual Acc.
MAESTRO Hawthorne et al. (2019) Curated (Competition recs.) Audio 2 Automatic F1-score
MusicNet Thickstun et al. (2017b) Curated (Classical recs.) Audio N/A Manual Recall
NES-MDB Donahue et al. (2018) Curated (Game Audio) Audio 11 Automatic N/A
MusicScore Lin et al. (2024) Curated (Public Scores) Score image corpus 8 Automatic FID
Lakh MIDI Raffel (2016) In-the-wild (Web-MIDI) MIDI N/A Automatic N/A
@@ 2.1 Symbolic Music Understanding and Benchmarks
Symbolic music understanding has traditionally been evaluated using clean, structured datasets such as MusicNet (Thickstun et al., 2017a), NES-MDB (Donahue et al., 2018), and MAESTRO (Hawthorne et al., 2019). These datasets align audio with symbolic formats to facilitate tasks like transcription and generation. However, they reflect highly curated environments, lacking the variability, ambiguity, and informal nature of user-generated content. Other symbolic corpora like MusicScore Lin et al. (2024) or Lakh MIDI Raffel (2016) further extend coverage but remain either score-centric or MIDI-based without real-world contextual grounding. Recent efforts like MusicTheoryBench Czajka et al. (2024b) introduce theory-centric evaluations, but they rely on expert-curated questions in controlled settings. WildScore differs by grounding symbolic music analysis in online discourse, incorporating informal reasoning and context-dependent ambiguity from platforms like Reddit Reddit (2024).
@@ 3 WildScore
The aim of this study is to evaluate the visual context understanding of Multimodal Large Language Models (MLLMs) for symbolic musical score as shown in Figure 2. To this end, we introduce WildScore. In this section, we describe the dataset details; Section 4 presents the evaluation of vision–language reasoning over symbolic musical scores.
Our dataset creation process involves two distinct phases: (1) data collection (3.1), (2) multimodal filtering (3.2). Together, these phases enforce sample relevance, symbol-image grounding, and rigorous quality control, yielding a benchmark that robustly evaluates MLLMs’ visual context understanding of symbolic music.
@@ 3.3 Dataset Categorization
To support structured analysis and evaluation, we categorized our dataset into five categories as shown in Figure 3 to represent core aspects of music theory. These categories are further divided into twelve detailed subcategories as shown in Figure  4:
@@ 3.4 Dataset Overview & Statistics
We assign each question to Easy, Medium, or Hard using an LLM-based rubric. Specifically, we prompt GPT-4.1 with a few-shot template designed from examples curated by a Level 5 expert (criteria in Table 6) to rate the expected difficulty from the prompt. The resulting distribution is as shown in Table 7. Additional construction details, annotator instructions, and prompt templates are provided in the appendix C.
@@ 4.2 Quantitative Results
GPT-4.1-mini OpenAI (2023) Params: undisclosed w/ Image 72.12 69.57 70.14 63.20 64.15 68.31
@@
Qwen-2.5-VL Bai et al. (2023) Params: 8.29B w/ Image 52.88 52.17 47.06 47.20 58.49 49.73
@@
Phi-3-Vision Abdin et al. (2024) Params: 4.15B w/ Image 45.19 52.17 48.42 48.00 56.60 48.82
@@
Gemma-3 Team et al. (2025) Params: 4.3B w/ Image 40.27 52.24 47.89 43.55 53.42 46.34
@@
MiniCPM Hu et al. (2024) Params: 3.43B w/ Image 50.00 45.65 44.34 44.80 47.17 45.90
@@
InternVL Chen et al. (2023) Params: 9.14B w/ Image 46.15 36.96 36.65 37.60 43.40 39.34
@@
LLaVA Liu et al. (2023b) Params: 7.06B w/ Image 37.50 41.30 28.96 32.00 35.85 32.97
@@
Table 4: Per‐subcategory accuracy (%) by model and input modality, with subcategories grouped by category.
A subcategory analysis (Table 4) reveals heterogeneous image contributions across models. For GPT-4.1-mini, accuracies peak on Dynamics and Articulation (87.18%) and Modal Mixture (79.25%) and drop on Orchestral Texture (33.33%) and Contrapuntal Forms (40.00%). Other systems show attenuated or even negative image gains in multiple subcategories. We hypothesize that this heterogeneity reflects differences in multimodal pretraining and alignment in models with stronger vision–language objectives and instruction tuning appear better grounded in symbolic notation, whereas those trained primarily on natural-image corpora or with weaker visual adapters show limited benefit from images. A Level-5 human expert was evaluated on 100 proportionally sampled questions spanning easy, medium, and hard categories, achieving an overall average accuracy of 72%. Additionally, besides Contrapuntal Forms most model performance without the image is better than random guessing, highlighting that the naturalized data used to create WildScore may not fully require perception of the scores. This makes an interesting contrast to recent synthetically difficult benchmarks that force multimodal perception to succeed Zang et al. (2025), as such difficult benchmarks may not reflect the real distribution of questions, where perception may not always be necessary.
@@ 5 Conclusion
Empirical results indicate that while current state-of-the-art MLLMs exhibit substantial promise, particularly in tasks involving surface-level recognition and straightforward analysis, they continue to struggle significantly with deep symbolic abstraction, rhythmic complexity, and orchestration intricacies especially when presented as an image. Significant differences in performance demonstrated by popular multimodal large language models between text-only and visual inputs highlight how important visual context is for precise musicological interpretation.
@@ Acknowledgments
This work was partially supported by the U.S. National Science Foundation under Grant IIS-2432486.